# Data-Adaptive Graph Framelets with Generalized Vanishing Moments for Graph Signal Processing

### v2

@@ V-C Denoising
[…]
VI Conclusion and Discussion
Our proposed method of constructing tight graph framelet systems is parameterized, general, and flexible in adjusting the space of scaling functions, resulting in locally supported framelets associated with fast forward and backward transforms and possessing generalized vanishing moments.
By learning the space $\mathbb{V}_{0}$ to capture the low-dimensional structure of a family of signals and learning the optimal bases/frames of the subspaces orthogonal to $\mathbb{V}_{0}$, our method can provide far more sparser representations compared with other non-data-adapted approaches. Such sparseness is reflected in the superior performance of non-linear approximation and denoising tasks.
The training time in solving the Riemannian optimization problems could be an issue. We would like to point out that some of the filters $\bm{A}^{[j,k]}$ can be predefined to reduce the number of parameters and, consequently, the training overhead. In this case, the system has a smaller representation capacity.
As for further applications, given a pre-trained system, the fast forward and backward transforms resemble the computation in general neural networks and can be combined with a properly defined neural network. Using its sparseness, a pre-trained system can serve as a regularizer in deep learning tasks. It can also applied in compressed sensing. These remain future directions to be explored.
Finally, our method can be applied to more general discrete and irregular data as long as a series of partitions is given. Thus, apart from graph data, it has the potential to be applied in more complicated settings.

### v3

@@ 6.2 Experiments
7 Conclusion and Discussion
We proposed a general system to generate frames on graphs that provides a practical solution to the following problem: sparsely representing arbitrary families of graph signals that are approximately low-dimensional. Our system is particularly suitable for the problem, as we proposed to learn the filters that adapt to the given signals. Apart from graph signal processing, we have shown that the system is also capable of generating rich and novel structural information that is beneficial to node classifications. However, the approach of selecting framelets for node classification is overall heuristic. In practice, this requires ad hoc and possibly a large number of attempts in choosing the framelets. It would be interesting to investigate whether there is any theoretical discussion or even guarantees for our framelet systems, as well as the ways of choosing framelets.
The training time consumed in the Riemannian optimizations could be an issue. We would like to point out that some of the filters $\bm{A}^{[j,k]}$ can be predefined to reduce the number of parameters and, consequently, the training overhead. In this case, the system has a smaller representation capacity. Specifically, a detailed investigation of the computational scalability of the proposed framelet systems would be necessary for big-data scenarios. As it is proposed to conduct certain optimizations on Stiefel manifolds in a specific manner, an analysis concerning the choice of algorithms is beneficial, and it is also interesting to consider exploiting the iteration patterns and backpropagation in GPU computing, see e.g. [67].
As for further applications, given a pre-trained system, the fast forward and backward transforms resemble the computation in general neural networks and can be combined with a properly defined neural network. Using its sparseness, a pre-trained system can serve as a regularizer in deep learning tasks. It can also be applied in compressed sensing. These remain future directions to be explored.
Finally, our method can be applied to more general discrete and irregular data as long as a series of partitions is given. Thus, apart from graph data, it has the potential to be applied in other settings.